# Making Every Frame Matter: Continuous Video Understanding for Large Models via Adaptive State Modeling

### v1

@@ 3.3. Temporal Relation-guided Spatial Redundancy Reduction
However, it is challenging, as the types of events in the real world are diverse. We cannot merely retain an object’s pose and position. For instance, in Figure 6, it is crucial to also keep the feature of balance beam present in the environment. In this way, we can accurately comprehend that this is a gymnastics action, rather than a simple jump.
To achieve the reduction goal, we first use a pre-trained powerful feature extractor to extract general-purpose spatial feature, and then use a lightweight projection layer to perform the reduction. Specifically, we utilize the CLIP (radford2021learning, ) to perform the preliminary spatial feature extraction. CLIP is one of the most powerful vision encoders, which serves as the vision encoder in mainstream video foundation models (awais2023foundational, ; madan2024foundation, ).
Then we utilize a linear layer followed by an activation function as the projection layer to reduce the redundancy all-purpose spatial features. The linear layer is trained together with our history modeling layer to reduce the redundancy with the guidance of the temporal relationship. After the reduction, the spatial feature can be used to enable efficient long-range history modeling, as shown on the far right of Figure 6. Thus, the complexity of modeling the relationship between two frames is reduced to $\mathcal{O}(1)$.
3.4. Parallel Training with Frame-weighted Loss
Given the recurrent form of our C-VUE may significantly affect the system performance during the training phase. We adopt a parallel training structure to efficiently train long video samples. Additionally, we design a frame-weighted loss to calculate the loss for each frame, thereby accomplishing prompt video understanding.
3.4.1. Efficient Parallel Training
To implement the redundancy reduction design described in Section 3.3, we start by using a pre-trained vision encoder ($\theta_{\texttt{spatio\_encoding}}$) to extract the general-purpose spatial feature. It is important to note that this vision encoder remains frozen during training. Before performing history modeling, we project the frame embeddings using a linear layer with ReLU activation ($\theta_{\texttt{reduction}}$). Recall that this projection is used to perform spatial redundancy reduction based on the temporal relationship. The $\theta_{\texttt{history\_modeling}}$ is the design proposed in Section 3.2 for adaptive history modeling. We unfold the recurrent state transition process into a convolutional form for parallel computation. Finally, we employ another linear layer with ReLU activation ($\theta_{\texttt{post\_proj}}$) to map the output of each time step to the space of the downstream task. The overall structure of our model is illustrated on the right side of Figure 7.
The training process is detailed in Algorithm 1. Although our system (C-VUE) employs a recurrent design, processing one frame at a time. We achieve high parallelism during training. Specifically, in the vision encoder, as well as during spatial feature reduction and post processing, we treat the temporal dimension as a batch dimension for parallel computation.
@@ 3.4.1. Efficient Parallel Training
             $\texttt{Emb}_{1:n}\leftarrow\theta_{\texttt{spatio\_encoding}}^{\star}(\texttt{X}_{1:n})$ // parallel
             $\texttt{Tok}_{1:n}\leftarrow\theta_{\texttt{reduction}}(\texttt{Emb}_{1:n})$ // parallel
             $\texttt{Ir}_{1:n}\leftarrow\theta_{\texttt{history\_modeling}}(\texttt{Tok}_{1:n})$ // in convolution mode
             $\hat{\texttt{Y}}_{1:n}\leftarrow\theta_{\texttt{post\_proj}}(\texttt{Ir}_{1:n})$ // parallel
@@ 5             $\theta\leftarrow\theta-\nabla_{\theta}\texttt{L}$
Algorithm 1 The parallel training structure of C-VUE. $\theta_{\texttt{spatio\_encoding}}^{\star}$ is frozen during the training.
Figure 7. The model architecture of our C-VUE.
3.4.2. Frame-weighted Loss for Continuous Understanding
Here we introduce the loss function $\mathcal{L}$ in Algorithm 1. Video datasets are often categorized as either motion-focused or scene-focused ones. The former requires understanding the process of action transitions to interpret the video accurately. For instance, in gymnastics sports datasets (shao2020finegym, ), the initial frames of turning once and twice may look similar, but the later frames are crucial for determining the action category. In contrast, scene-focused tasks can often be distinguished by certain static appearances, such as drinking water or reading (sigurdsson2016hollywood, ).
Video datasets are usually annotated in a temporal-grounded manner. To account for the different characteristics of the video dataset, we propose a frame-weighted loss. The key principle is to compute the loss by taking a weighted average across all frames. Formally, this can be expressed as:
(7) $\mathcal{L}_{v}=W^{T}\times[\mathcal{L}_{f}^{1},\mathcal{L}_{f}^{2},\ldots,\mathcal{L}_{f}^{N}],$
where $W$ is the weight vector of the loss from Frame 1 to Frame $N$.
@@ 3.4.2. Frame-weighted Loss for Continuous Understanding
Users can design different weight vectors based on their datasets. For motion-focused datasets, we recommend a linear growth strategy, where the weight of the loss for the initial frame is relatively low and gradually increases over time within the video frames. Specifically, $W=[\frac{i}{N^{2}}]_{i=1}^{N}$. Conversely, for scene-focused datasets, a uniformly weighted approach is recommended. That is, $W=[\frac{1}{N}]_{i=1}^{N}$.
4. Implementation
Network Architecture. We implement the $\theta_{\texttt{spatio\_encoding}}^{\star}$ with the powerful vision encoder, i.e., CLIP (clip_impl, ). $\theta_{\texttt{temporal\_modeling}}$ is implemented with a single layer Mamba (gu2023mamba, ). Here we utilize Mamba implementation for acceleration. Mamba proposes a selective scanning algorithm, which assumes the order in which it does operations does not matter through the associate property. It calculates the sequences in parts and iteratively combines them to calculate the history modeling in a convolution mode.
@@ 4. Implementation
Hyperparameter Setting. During the training, we freeze the vision encoder and only train the two projection layers and the Mamba layer. The AdamW (loshchilov2017decoupled, ) is adopted as the optimizer. The learning rate changes dynamically, using Cosine annealing with warm restarts. This method adjusts the learning rate between 1e-3 and 1e-6 every 10 epochs. Aside from these adjustments, the other hyperparameters are kept at their default settings as provided in PyTorch v2.2.0. Regarding the training process of the network, it requires approximately 100 to 200 epochs to converge in different datasets. We train these models on a server with 8 RTX4090 GPU cards, and it takes about 3 to 6 hours to train a model on different datasets.
Optimization Deployment. To ensure real-time operation and cross-platform deployment, we have utilized compilation and quantization techniques on our model. We have encoded the model file into $C$ code to guarantee cross-platform deployment, and ultimately achieved deployment on a MacBook Air (M3, 2024), Jetson AGX Orin 64 GB, and a PC with an RTX 4090 GPU card.
@@
We implemented our vision encoder CLIP (radford2021learning, ) based on the llama.cpp framework (llama_cpp_impl, ), completed the compilation, and carried out INT8 quantization. The existing CLIP.cpp (clip_cpp_impl, ) framework based on llama.cpp for compiling CLIP only supports CPU backend, hence it can’t utilize the GPU on mobile devices. Other mainstream frameworks lack good compatibility for cross-platform deployment. For example, OpenClip (open_clip_impl, ) and ONNXRuntime-GPU (onnx_impl, ) lack support for using the CLIP model with Jetson GPU. We have expanded support for ggml in our implementation, enabling GPU-based inference on all the aforementioned platforms, which resulted in significant system performance improvement. Please refer to Figure 9. The official implementation of Mamba (mamba_impl, ) only supports the PyTorch framework with the CUDA backend. We have compiled our Mamba part into a $C$ program, thereby allowing its deployment on various platforms.
5. Evaluation
5.1. Experiment Setups
Evaluation Goals. We aim to evaluate five capabilities of the system:
Capability 1: Understanding of events with various lengths.
Capability 2: Prompt response as more frames are received.
Capability 3: Generality for long-tail data.
Capability 4: Understanding of continuous input events.
Capability 5: Long-range memorization.
@@ 5.1. Experiment Setups
Figure 8. Datasets used in experiments. S represents a sample; a line segment represents the duration of an event; E represents a label. In Charades and Multi-label FineGym, a frame may have multiple labels.
@@
Datasets. We utilize five datasets to evaluate different capabilities. FineGym99 (shao2020finegym, ) (For capabilities 1 & 2). It is a gymnasium video dataset. Each video clip is labeled with a gymnastics movement. The duration of events varies from 2 to 50 seconds.
FineGym288 (shao2020finegym, ) (For capabilities 1, 2, & 3). It is an extended version of FineGym99, which enhances its by adding categories. FineGym288 has the natural long-tail distribution.
@@
Charades (sigurdsson2016hollywood, ) (For capabilities 1, 2, & 5). It is a multi-label dataset, composed of daily indoor activities with an average length of 30 seconds. Each video clip consists of about 800 frames on average, varying from 32 to 3,944 frames. We have re-annotated the video samples for continuous video understanding as Figure 8(b).
@@
All-event-in-one FineGym (For capabilities 1, 2, & 4). It is constructed based on the FineGym99 as shown in Figure 8(c). We concatenated all the video clips from FineGym99’s testing dataset, forming a single long video lasting 3.5 hours and consisting of over 440k frames.
Multi-Label FineGym@k (For capabilities 1, 2, & 5). It is constructed based on the FineGym99 as shown in Figure 8(d). The new video clips are constructed by randomly concatenating $k$ events of FineGym99 together and re-annotating them.
@@
Baseline Methods. We utilize five baselines to execute continuous perception, whereby the method predicts an outcome at each timestep. Baselines 1, 2, and 3 are constructed following different video processing paradigms. Baselines 4 and 5 are state-of-the-art (SOTA) methods for balancing accuracy and efficiency for video classification.
Baseline 1: Transformer (vaswani2017attention, ). The overall architecture mirrors ours for a fair comparison, except that we replace our history modeling layer with a transformer of equivalent parameter size.
Baseline 2: ConvPooling (kar2017adascan, ). The overall architecture mirrors ours for a fair comparison, except that we replace the history modeling layer with a convPooling of equivalent parameter size. The convPooling consists of a 1D convolution and a 1D adaptive max pooling.
Baseline 3: LSTM (hochreiter1997long, ). This design keeps the overall structure consistent with ours for a fair comparison, substituting the history modeling layer with an LSTM of equal parameter magnitude.
Baseline 4: MoviNetA0 (kondratyuk2021movinets, ). This is a Mobile Video Network specifically designed for efficient video recognition. This method applies down-sampling when processing videos. Drawing from the experience of using this model on the Charades dataset, we similarly down-sample the video to 6 fps for video preprocessing.
@@
Baseline 5: Mvitv2(li2022mvitv2, ) This is an improved multiscale vision transformer for both classification and detection tasks. This method necessitates the sampling of any given video to a fixed number of frames for video classification tasks. We trained the baseline that takes as input 16 frames at a time, which is one of the default settings.
@@
S1: All frames feeding. All frames from frame 1 up to frame $N$ are fed into the baseline method to provide the necessary context for achieving continuous perception at timestep $N$.
S2: Sliding windowed feeding. All frames between the $(N-m)$-th frame to the $N$-th frame are fed into the baseline at timestep $N$, where $m$ denotes the sliding window size.
S3: Single frame feeding. Only frame $N$ is supplied to the baseline method.
Metrics. In this work, we utilize various metrics, namely Accuracy, Mean Average Precision (mAP), and the Jaccard Index for single-label and multi-label video classification tasks. Furthermore, we introduce a novel metric referred to as the Early Detection Rate (EDR), designed to evaluate the model’s ability to detect events promptly within videos. The definition is $\mathtt{EDR}={(E_{pred}-E_{gt})}/{E_{N}}$, where $E_{pred}$ refers to the frame index at which the label is detected. $E_{gt}$ represents the ground truth frame index, that is, the index of the frame where the event begins. $E_{N}$ signifies the total number of frames in the event.
Table 1. The accuracy and EDR results of different baselines. The #Param counts the parameters that need to be trained. Baselines 1, 2, 3, and our method have the same pre-trained clip with a parameter size of 149M.
@@
Baseline 1 (Transformer) 210.5K 82.3 48.20 76.6 53.20 17.2 21.4 52
Baseline 2 (ConvPooling) 367.3K 82.6 38.40 75.4 41.40 15.8 21 43.3
Baseline 3 (LSTM) 218K 84.1 23.40 78.4 25.80 15.33 17.8 84
Baseline 4 (MoviNet) 2.1M 19.7 41.60 13.9 44.20 3.9 5.8 32
Baseline 5 (Mvitv2) 34.3M 6.4 20.50 3.1 24.10 5.3 7.3 99.9

### v2

@@ 3.1 Activity Spatial Feature Extraction
The challege of addressing issue 1 is that we lack supervised data to differentiate between redundant spatial features and those relevant to activities. Because the types of activities in the real world are diverse. We cannot merely retain an object’s pose and position. For instance, in Figure 5, it is crucial to also keep the feature of the balance beam present in the environment. In this way, we can accurately comprehend that this is a gymnastics activity, rather than a simple jump.
The core idea of our feature extraction is to utilize the information bottleneck [38] as an unsupervised method to extract task-specific features. Specifically, we employ an information reduction module aimed at minimizing input features while maximizing task accuracy during training. Through this adversarial constraint, it can reduce redundancy and effectively achieve activity-related feature extraction.
Formally speaking, let $X$, $X^{\prime}$, and $Z$ denote the random variables that represent all spatial features, extracted activity-related features, and redundant features, respectively. In our scenario, the activity spatial feature extraction determines how the feature $X^{\prime}$ is extracted from $X$ for the recognition task $f_{\theta}$. The goal of the activity-related feature extraction is to minimize the inclusion of redundancy information $Z$ in extracted feature $X^{\prime}$ while ensuring maximum data utility, i.e.,
$\min_{\mathcal{I}(X^{\prime};Z)<\epsilon}\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|,$ (1)
[…]
However, given the lack of supervised data regarding redundant information $Z$, it is hard to minimize the $\mathcal{I}(X^{\prime};Z)$. Note that, $\mathcal{I}(X^{\prime};Z)\leq\mathcal{H}(X^{\prime})$, where $\mathcal{H}$ is the entropy of a random variable. So, we have the relaxation of Equation 2, i.e.,
$\min_{\mathcal{H}(X^{\prime})<\epsilon}\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|.$ (2)
That is, by minimizing the information of extracted features $X^{\prime}$ and preserving the accuracy of activity recognition, we can remove the redundant features as much as possible. Please note that the data utility objective, $\min\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|$, is optimized by the accuracy of the activity recognition task (Equation 5).
To reduce the information entropy of extracted features $X^{\prime}$, we propose two steps, i.e., the pooling layer and an information funnel Specifically, in our activity spatial feature extraction module, we first a pre-trained powerful feature extractor, i.e., CLIP [29], to extract general-purpose spatial features from an input frame. The CLIP model divides the input frame into multiple patches and computes an embedding for each patch. We utilize the average pooling to aggregate these embeddings by computing their mean, generating a single embedding. This constitutes the first stage of information reduction for $X^{\prime}$. After the first stage, each frame is reduced to a single embedding. Subsequently, this embedding is passed through our information funnel, which consists of a linear layer followed by an activation function, where the output size of the linear layer is smaller than its input size. This completes the second stage of information reduction for $X^{\prime}$.
3.2 Activity-Aware State Update
The challege of addressing issue 2 lies in online associating the features of each incoming frame with those in the video context, i.e., hidden state, thereby enhancing recognition accuracy. This state update serves as the association process since the recognition results are based on the updated state.
We take the RNN as an example to introduce the traditionalstate update. It computes $\mathbf{y}$ through a dynamic recurrence of input signals $\mathbf{x}$ at each time step $t$ through hidden states $\mathbf{s}$. Formally,
$s_{t}$ $=\phi_{h}(\mathbf{A}s_{t-1}+\mathbf{B}x_{t}+\mathbf{b}_{h}),$ (3)
$y_{t}$ $=\phi_{o}(\mathbf{C}s_{t}+\mathbf{b}_{o}).$
where $\mathbf{A},\mathbf{B}$, and $\mathbf{C}$ denote the state matrix, input matrix, and output matrix, respectively. The $\mathbf{b}$ denotes the bias parameters, and the functions $\phi$ denotes activation functions.
However, the RNN and its gating-based variants, e.g., LSTM [31], are widely considered challenging to maintain input context by many research works [12, 45, 27]. Our experimental results indicate that the LSTM-based baseline performs poorly on streaming video inputs (Table 2). We argue that the underperformance of such methods is due to their inefficient utilization of the hidden state. Specifically, standard RNN processes input at different moments identically, i.e., fixed $A$, $B$, and $C$ in Equations 3, leading to irrational use of the hidden state (Figure 2).
We argue that this deficiency stems from the state update method’s inability to adapt effectively to incoming video frames, as it applies one update strategy to all frames, irrespective of their distinct features. We enhance the state update method by enabling activity-aware hidden state updates based on activity spatial features. The core idea is to introduce dynamic adaptability, akin to the attention mechanism in transformers, which adjusts its focus across input sequences to accommodate varying inputs.
Specifically, we replace the fixed matrix used for state updates with an input-dependent update manner. Our input-dependent update takes the current activity spatial features as input and determines how the current input should be incorporated into the hidden state. Formally,
$A\rightarrow A_{t}=f_{A}(x_{t});B\rightarrow B_{t}=f_{B}(x_{t});C\rightarrow C_{t}=f_{C}(x_{t}).$ (4)
That is, during training, we no longer learn the $A$, $B$, and $C$ matrices directly but instead learn how to dynamically generate these matrices based on the input.
Here, we employ Mamba [13] as the implementation of our activity-aware state update module, as it aligns with our design philosophy: it dynamically computes these matrices based on each input. Additionally, Mamba’s implementation accelerates the training process for such recursive formulations. Note that while Mamba has been explored for video processing [20, 8, 7], these works adopt strategies such as processing “all frames at once” or using a “fixed window size”, treating Mamba primarily as a lightweight alternative to Transformers. In contrast, our “state-assistant” design fundamentally differs in both approach and purpose.
3.3 Training Structure
The training structure is detailed in Algorithm 1. The CARS’s architecture consists of a pre-trained vision encoder ($\theta_{\texttt{CLIP}}$), a feature extractor ($\theta_{\texttt{extractor}}$), an activity-aware state update module ($\theta_{\texttt{state\_maintaince}}$), and a classifier ($\theta_{\texttt{cls}}$). The vision encoder extracts the general-purpose spatial feature and is frozen during training. The feature extractor is a linear layer with ReLU activation. The activity-aware state update module is a single Mamba layer. The classifier is a ReLU activation.
@@ 3.3 Training Structure
4             $\texttt{Emb}_{1:n}\leftarrow\theta_{\texttt{CLIP}}^{\star}(\texttt{X}_{1:n})$
5             $\texttt{Tok}_{1:n}\leftarrow\theta_{\texttt{extractor}}(\texttt{Emb}_{1:n})$
6             $\texttt{Ir}_{1:n}\leftarrow\theta_{\texttt{state\_maintaince}}(\texttt{Tok}_{1:n})$
7             $\hat{\texttt{Y}}_{1:n}\leftarrow\theta_{\texttt{cls}}(\texttt{Ir}_{1:n})$
@@ 9             $\theta\leftarrow\theta-\nabla_{\theta}\texttt{L}$
Algorithm 1 The parallel training structure of CARS. $\theta_{\texttt{spatio\_encoding}}^{\star}$ is frozen during the training.
The loss function $\mathcal{L}$ is a frame-weighted loss. The core intuition is that activity recognition can be categorized into two types: motion-focused (requiring temporal information, e.g., performing a triple somersault) or scene-focused (relying solely on static information, e.g., drinking water). Formally, it can be expressed as:
$\mathcal{L}_{v}=W^{T}\times[\mathcal{L}_{f}^{1},\mathcal{L}_{f}^{2},\ldots,\mathcal{L}_{f}^{N}],$ (5)
@@
where $W$ is the weight vector of the loss from Frame 1 to Frame $N$. For motion-focused tasks, we recommend a linear growth strategy, where the weight of the loss for the initial frame is relatively low and gradually increases over time within the video frames. Specifically, $W=[\frac{i}{N^{2}}]_{i=1}^{N}$. Conversely, for scene-focused datasets, a uniformly weighted approach is recommended. That is, $W=[\frac{1}{N}]_{i=1}^{N}$.
4 Evaluation
@@
Hyperparameter Setting. The AdamW [24] is adopted as the optimizer. The learning rate changes dynamically, using Cosine annealing with warm restarts. This method adjusts the learning rate between 1e-3 and 1e-6 every 10 epochs. Aside from these adjustments, the other hyperparameters are kept at their default settings as provided in PyTorch v2.2.0. Regarding the training process of the network, it requires approximately 100 to 200 epochs to converge in different datasets. We train these models on a server with 8 RTX4090 GPU cards, and it takes about 3 to 6 hours to train a model on different datasets.
@@
Optimization Deployment. We implemented the CLIP vision encoder using the llama.cpp framework [2], performing INT8 quantization for improved efficiency. While the CLIP.cpp framework [1], based on llama.cpp, only supports CPU backends, limiting GPU use on mobile devices, other frameworks like OpenClip [5] and ONNXRuntime-GPU [4] also lack support for Jetson GPUs. We extended support for ggml format, enabling GPU-based inference across all platforms, which resulted in a significant performance boost (see Figure 9). Furthermore, we compiled the Mamba implementation [3], which originally supported only PyTorch with CUDA, into ggml format for cross-platform deployment. Through these optimizations, we successfully deployed our CARS on a MacBook Air (M3, 2024), Jetson AGX Orin 64GB, and a PC with an RTX 4090 GPU.
@@
Figure 7: Datasets used in experiments. S represents a sample; a line segment represents the duration of an activity; E represents a label. In Charades, a frame may have multiple labels.
4.1 Experiment Setups
Evaluation Goals. We aim to evaluate four capabilities:
Capability 1: Recognition of activities across timescales.
Capability 2: Activity Recognition in streaming input.
Capability 3: Prompt response as more frames are received.
Capability 4: Generality for long-tail activities.
@@ 4.1 Experiment Setups
Datasets. We utilize five datasets to evaluate different capabilities. FineGym99 [30] (For capabilities 1 & 3). It is a gymnasium video dataset. Each video clip is labeled with a gymnastics movement. The duration of activities varies from 2 to 50 seconds. FineGym288 [30] (For capabilities 1, 3, & 4). It is an extended version of FineGym99, which enhances its by adding categories. FineGym288 has a natural long-tail distribution.
@@
Charades [34] (For capabilities 1 & 3). It is a multi-label dataset, composed of daily indoor activities with an average length of 30 seconds. Each video clip consists of about 800 frames on average, varying from 32 to 3,944 frames. We re-annotated Charades, as Figure 7(b), to evaluate the method’s effectiveness in maintaining video context when multiple activities occur simultaneously.
@@
All-activity-in-one FineGym (For capabilities 1, 2, & 3). It is constructed based on the FineGym99 as shown in Figure 7(c). We concatenated all the video clips from FineGym99’s testing dataset, forming a single long video lasting 3.5 hours and consisting of over 440k frames. The concatenated video is employed to simulate streaming video containing continuous activities.
@@
Baseline Methods. We utilize five baselines to execute continuous perception, whereby the method predicts an outcome at each timestep. Baselines 1, 2, and 3 are constructed following different video processing paradigms (as shown in Figure 8). Their overall architecture mirrors ours, except that we replace the mamba layer with other networks. Baseline 1 utilizes a Transformer [39]. Baseline 2 utilizes a convPooling [18], which consists of a 1D convolution and a 1D adaptive max pooling. Baseline 3 utilizes an LSTM [14].
@@
Baselines 4 and 5 are state-of-the-art (SOTA) methods for balancing accuracy and efficiency for video classification for on-device computing scenarios. Baseline 4: MoviNetA0 [19]. This is a Mobile Video Network specifically designed for efficient video recognition. This method applies down-sampling when processing videos. Drawing from the experience of using this model on the Charades dataset, we similarly down-sample the video to 6 fps for video preprocessing. Baseline 5: Mvitv2[22] This is an improved multiscale vision transformer for both classification and detection tasks. This method necessitates the sampling of any given video to a fixed number of frames for video classification tasks. We trained the baseline that takes as input 16 frames at a time, which is one of the default settings.
Figure 8: The constructed baselines 1, 2, and 3.
@@
S1: All frames feeding. All frames from frame 1 up to frame $N$ are fed into the baseline method to provide the necessary context for achieving continuous perception at timestep $N$. S2: Sliding windowed feeding. All frames between the $(N-m)$-th frame to the $N$-th frame are fed into the baseline at timestep $N$, where $m$ denotes the sliding window size. S3: Single frame feeding. Only frame $N$ is supplied to the baseline method.
Metrics. In this work, we utilize various metrics, namely Accuracy, Mean Average Precision (mAP), and the Jaccard Index for single-label and multi-label video classification tasks. Furthermore, we introduce a novel metric referred to as the Early Detection Rate (EDR), designed to evaluate the model’s ability to detect activities promptly within videos. The definition is $\mathtt{EDR}={(E_{pred}-E_{gt})}/{E_{N}}$, where $E_{pred}$ refers to the frame index at which the label is detected. $E_{gt}$ represents the ground truth frame index, that is, the index of the frame where the activity begins. $E_{N}$ signifies the total number of frames in the activity.
Table 1: The accuracy and EDR results of different baselines. The #Param counts the parameters that need to be trained. Baselines 1, 2, 3, and our method have the same pre-trained clip with a parameter size of 149M.
@@
B1 (Transformer-based) 210.5K 82.3 48.20 76.6 53.20 17.2 21.4 52
B2 (ConvPooling-based) 367.3K 82.6 38.40 75.4 41.40 15.8 21 43.3
B3 (LSTM-based) 218K 84.1 23.40 78.4 25.80 15.33 17.8 84
B4 (MoviNet) 2.1M 19.7 41.60 13.9 44.20 3.9 5.8 32
B5 (Mvitv2) 34.3M 6.4 20.50 3.1 24.10 5.3 7.3 99.9